According to Bohrs theory of hydrogen atom angular momentum of electron in any orbit is directly proportional to square root of the radius of orbit. HOW? 

Grade:upto college level

4 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student

Angular momentum = mvr
v=root(ZKe^2/mr)
So Angular Momentum = mvr=root(ZKe^2*r*m)
and that how its proportional to square root of the radius

Thanks & Regards
Arun Kumar
Askiitians Faculty
IIT Delhi
Ashish Sahani
6 years ago
We know that r=n^2/Z. ............(1)For Hydrogen, Z=1SoEquation (1) became r=n^2orn=√r. ..................(2)We also know thatAngular momentum =nh/2π. ......(3)From eq. (2)We can writeAngular momentum is directly proportional to √r
Ashish Sahani
6 years ago
We know that r=n^2/Z. ............(1)For Hydrogen, Z=1SoEquation (1) became r=n^2orn=√r. ..................(2)We also know thatAngular momentum =nh/2π. ......(3)From eq. (2) [n =√r]We can writeAngular momentum is directly proportional to √r
Kushagra Madhukar
3 years ago
Dear student,
Please find the answer to your question.
 
We know that angular momentum = r x linear momentum = r x P
= r x 2m(KE)
= r x 2m(KZe2/2r)
= KZe2mr
Hence angular momentum, L ∝ r
 
Thanks and regards,
Kushagra